# Is Mamba Capable of In-Context Learning?

### v1

@@ A.2.6 ICL learning curves
Figure 8: Estimated scalar scale and shift by the linear probing model used in Section 2.2.
As mentioned above, for constructing the linear model, we selectively used the representation from the in-context examples ranging from the 30th to the 40th. This selective usage was due to the observation that the model’s initial predictions are suboptimal, as evidenced by the performance comparison in Figure 1. By focusing on later samples, we aim to derive a more accurate linear model based on the embeddings corresponding to more refined predictions.
The scale and shift were estimated for each layer and are reported in Figure 8. We note that for skewed linear regression and ReLU neural networks the value for scale approaches 1, and the one for shift approaches 0 relatively quickly.
A.3 ICL for NLP tasks
Table 1: Complete results of the ICL benchmark by Hendel et al. (2023) for all tasks and models. Regular denotes the ICL setting as described in Section 3 and baseline refers to zero-shot performance (i.e. a query without any examples).
@@ A.3 ICL for NLP tasks
Next letter 0.16 0.76
Prev letter 0.02 0.02
@@
Knowledge Country capital 0.08 0.20
Location continent 0.00 0.55
Location religion 0.02 0.67
Person language 0.02 0.48
Linguistic Antonyms 0.19 0.30
Plural singular 0.06 0.44
Present simple gerund 0.00 0.67
Present simple past simple 0.02 0.69
Translation En es 0.06 0.25
En fr 0.14 0.37
Es en 0.09 0.21
Fr en 0.08 0.25
@@
Next letter 0.04 0.87
Prev letter 0.09 0.12
@@
Knowledge Country capital 0.10 0.53
Location continent 0.01 0.69
Location religion 0.02 0.59
Person language 0.02 0.75
Linguistic Antonyms 0.29 0.67
Plural singular 0.04 0.74
Present simple gerund 0.00 0.86
Present simple past simple 0.00 0.81
@@
Next letter 0.09 0.88
Prev letter 0.04 0.34
@@
Knowledge Country capital 0.18 0.71
Location continent 0.03 0.73
Location religion 0.12 0.73
Person language 0.01 0.79
Linguistic Antonyms 0.39 0.76
Plural singular 0.12 0.86
Present simple gerund 0.00 0.91
Present simple past simple 0.00 0.85
Translation En es 0.20 0.56
En fr 0.16 0.58
Es en 0.18 0.78
Fr en 0.13 0.75
@@
Next letter 0.07 0.81
Prev letter 0.09 0.34
@@
Knowledge Country capital 0.20 0.71
Location continent 0.02 0.76
Location religion 0.04 0.83
Person language 0.00 0.83
Linguistic Antonyms 0.30 0.77
Plural singular 0.11 0.86
Present simple gerund 0.00 0.91
Present simple past simple 0.00 0.92
Translation En es 0.17 0.73
En fr 0.24 0.67
Es en 0.07 0.83
Fr en 0.11 0.78
Mamba 2.80B Algorithmic List first 0.69 1.00
@@
Next letter 0.05 0.93
Prev letter 0.00 0.39
@@
Knowledge Country capital 0.20 0.75
Location continent 0.02 0.82
Location religion 0.07 0.82
Person language 0.00 0.86
Linguistic Antonyms 0.25 0.77
Plural singular 0.17 0.90
Present simple gerund 0.00 0.95
Present simple past simple 0.00 0.93
Translation En es 0.11 0.78
En fr 0.12 0.69
Es en 0.05 0.83
Fr en 0.11 0.78
@@
Next letter 0.12 0.82
Prev letter 0.02 0.22
@@
Knowledge Country capital 0.31 0.81
Location continent 0.04 0.86
Location religion 0.07 0.87
Person language 0.03 0.83
Linguistic Antonyms 0.34 0.78
Plural singular 0.11 0.89
Present simple gerund 0.00 0.90
Present simple past simple 0.02 0.92
Translation En es 0.09 0.73
En fr 0.09 0.59
Es en 0.12 0.81
Fr en 0.11 0.80
@@
Next letter 0.12 0.27
Prev letter 0.04 0.03
To lower 0.00 0.57
To upper 0.00 0.86
Knowledge Country capital 0.14 0.09
Location continent 0.00 0.67
Location religion 0.03 0.62
Person language 0.00 0.33
Linguistic Antonyms 0.09 0.06
Plural singular 0.16 0.22
Present simple gerund 0.00 0.21
Present simple past simple 0.01 0.14
Translation En es 0.14 0.14
En fr 0.11 0.25
Es en 0.02 0.09
Fr en 0.14 0.09
@@
Next letter 0.07 0.64
Prev letter 0.00 0.02
@@
Knowledge Country capital 0.14 0.34
Location continent 0.00 0.58
Location religion 0.10 0.69
Person language 0.01 0.59
Linguistic Antonyms 0.10 0.58
Plural singular 0.13 0.21
Present simple gerund 0.00 0.53
Present simple past simple 0.01 0.59
Translation En es 0.04 0.38
En fr 0.03 0.31
Es en 0.10 0.58
Fr en 0.08 0.52
@@
Next letter 0.22 0.78
Prev letter 0.00 0.04
@@
Knowledge Country capital 0.12 0.40
Location continent 0.00 0.56
Location religion 0.02 0.70
Person language 0.00 0.76
Linguistic Antonyms 0.31 0.68
Plural singular 0.07 0.70
Present simple gerund 0.00 0.70
Present simple past simple 0.01 0.79
Translation En es 0.21 0.63
En fr 0.14 0.57
Es en 0.15 0.78
Fr en 0.08 0.77

### v2

@@ A.1.6  ICL learning curves
Figure 12: Estimated scale and shift by the linear probing model used in Section 3.2.
The scale and shift were estimated for each layer and each task and are reported in Figure 12. We note that for skewed linear regression and ReLU neural networks the value for scale approaches 1, and the one for shift approaches 0 relatively quickly.
A.2  ICL for NLP tasks
We added Mamba and RKWV to the code provided by (Hendel et al.,, 2023) at github.com/roeehendel/icl_task_vectors. We leveraged their results for GPT-J, Llama and Pythia and provided our experimental results for Mamba and RWKV. A detailed break down of the per task performance of each model is given in Table LABEL:table:nlp_icl_performance.
Table 1: Complete results of the main experiment for all tasks and models.
@@ A.2  ICL for NLP tasks
Next letter 0.24 0.76
Prev letter 0.06 0.04
@@
Knowledge Country capital 0.07 0.19
Location continent 0.01 0.50
Location religion 0.02 0.63
Person language 0.01 0.53
Linguistic Antonyms 0.26 0.34
Plural singular 0.09 0.46
Present simple gerund 0.00 0.68
Present simple past simple 0.01 0.63
Translation En es 0.07 0.23
En fr 0.16 0.39
Es en 0.07 0.26
Fr en 0.10 0.27
@@
Next letter 0.19 0.87
Prev letter 0.07 0.11
@@
Knowledge Country capital 0.10 0.57
Location continent 0.00 0.64
Location religion 0.06 0.56
Person language 0.01 0.79
Linguistic Antonyms 0.33 0.68
Plural singular 0.09 0.74
Present simple gerund 0.00 0.83
Present simple past simple 0.00 0.80
@@
Next letter 0.12 0.85
Prev letter 0.03 0.35
@@
Knowledge Country capital 0.15 0.76
Location continent 0.02 0.73
Location religion 0.12 0.74
Person language 0.00 0.83
Linguistic Antonyms 0.45 0.76
Plural singular 0.07 0.86
Present simple gerund 0.00 0.90
Present simple past simple 0.01 0.86
Translation En es 0.11 0.56
En fr 0.14 0.61
Es en 0.09 0.82
Fr en 0.20 0.75
@@
Next letter 0.07 0.77
Prev letter 0.05 0.41
@@
Knowledge Country capital 0.15 0.76
Location continent 0.01 0.77
Location religion 0.06 0.77
Person language 0.00 0.84
Linguistic Antonyms 0.37 0.78
Plural singular 0.07 0.87
Present simple gerund 0.00 0.89
Present simple past simple 0.01 0.87
Translation En es 0.09 0.72
En fr 0.14 0.70
Es en 0.11 0.82
Fr en 0.14 0.72
Mamba 2.80B Algorithmic List first 0.70 1.00
@@
Next letter 0.02 0.95
Prev letter 0.04 0.37
@@
Knowledge Country capital 0.15 0.80
Location continent 0.03 0.81
Location religion 0.12 0.79
Person language 0.02 0.89
Linguistic Antonyms 0.38 0.79
Plural singular 0.14 0.96
Present simple gerund 0.00 0.96
Present simple past simple 0.01 0.93
Translation En es 0.03 0.75
En fr 0.07 0.67
Es en 0.11 0.83
Fr en 0.16 0.81
@@
Next letter 0.17 0.84
Prev letter 0.02 0.23
@@
Knowledge Country capital 0.20 0.84
Location continent 0.06 0.86
Location religion 0.12 0.84
Person language 0.04 0.87
Linguistic Antonyms 0.36 0.77
Plural singular 0.11 0.90
Present simple gerund 0.00 0.91
Present simple past simple 0.01 0.90
Translation En es 0.08 0.74
En fr 0.15 0.62
Es en 0.11 0.83
Fr en 0.16 0.76
@@
Next letter 0.10 0.29
Prev letter 0.00 0.03
To lower 0.00 0.54
To upper 0.00 0.85
Knowledge Country capital 0.17 0.10
Location continent 0.02 0.67
Location religion 0.10 0.60
Person language 0.01 0.31
Linguistic Antonyms 0.10 0.08
Plural singular 0.12 0.24
Present simple gerund 0.00 0.25
Present simple past simple 0.00 0.20
Translation En es 0.04 0.18
En fr 0.11 0.24
Es en 0.08 0.10
Fr en 0.17 0.10
@@
Next letter 0.09 0.66
Prev letter 0.00 0.01
@@
Knowledge Country capital 0.10 0.37
Location continent 0.00 0.56
Location religion 0.07 0.71
Person language 0.01 0.64
Linguistic Antonyms 0.16 0.56
Plural singular 0.10 0.22
Present simple gerund 0.00 0.51
Present simple past simple 0.01 0.64
Translation En es 0.04 0.41
En fr 0.18 0.39
Es en 0.08 0.57
Fr en 0.08 0.49
@@
Next letter 0.20 0.81
Prev letter 0.00 0.05
@@
Knowledge Country capital 0.13 0.41
Location continent 0.02 0.58
Location religion 0.04 0.73
Person language 0.00 0.79
Linguistic Antonyms 0.37 0.70
Plural singular 0.03 0.69
Present simple gerund 0.00 0.71
Present simple past simple 0.00 0.74
Translation En es 0.11 0.60
En fr 0.15 0.61
Es en 0.13 0.82
Fr en 0.14 0.73
RWKV 3B Algorithmic List first 0.29 0.96
List last 0.43 0.88
Next letter 0.02 0.56
Prev letter 0.02 0.08
To lower 0.00 1.00
To upper 0.00 1.00
Knowledge Country capital 0.08 0.79
Location continent 0.00 0.74
Location religion 0.02 0.74
Person language 0.00 0.82
Linguistic Antonyms 0.36 0.74
Plural singular 0.09 0.82
Present simple gerund 0.00 0.83
Present simple past simple 0.01 0.83
Translation En es 0.07 0.61
En fr 0.13 0.63
Es en 0.01 0.84
Fr en 0.05 0.75